I am having a tough time with these problems on this one I must use the discriminate to determine the number of solutions of the quadratic equation and whether the solutions are real or complex. It is not necessary to find the roots just determine the number of types of solutions.
4/3x^2 - 2x + 3/4 =0

Answer by rfer(10417)

the discriminant
b^2-4ac
------------------
4/3x^2-2x+3/4=0
a=4/3, b=-2, c=3/4
------------------
-2^2-4(4/3)(3/4)
----------------
Does that help?
Bob